Compute the number of passwords of each type below along with how long it would taketo...

  1. Compute the number of passwords of each type below along with how long it would taketo test all possible such passwords if it takes 1 nanosecond to test a password. Report the times in the most convenient human understandable form. In all of these, unless noted otherwise, order matters and repetition of characters is allowed.
    1. Passwords of length 8 with any combination of lowercase letters, uppercase letters,and numbers.
    2. Passwords that start with a capital letter, have 10 lowercase letters, and end with a number.
    3. Passwords that start with a capital letter, have between 10 and 20 lowercase letters,and end with a number.
    4. Passwords that have four words in them, where the words come from a list of 20,000 words.
    5. Passwords using the word ‘password’ with different uses of upper case and lowercase letters, and also allowing for a substitution of a number for the letters such as ‘96553014’; each letter is uniquely represented by a number, e.g., p can be replaced with 9 but with no other number. Basically, we have 3 choices per letter in ‘password’. No reordering allowed.
    6. Which of these schemes would you use for your bank account?
Answer #1

(there are more than 4 parts so as per policy i will answer forst 4 : a,b,c,d)

a.

total options for each unit of the password = 26 lowercase letters + 26 uppercase letters + 10 numbers

= 62

password length = 8

no. of possible passwords = (no. of options per unit)^(length)

= 62^8

= 2.1834011 * 10^14

time to test = (no. of possible passwords) * (time to test each)

= ( 2.1834011 * 10^14 ) * (1 nanosecond)

= 2.1834011 * 10^14 nanoseconds

b.

no. of passwords = (no. of uppercase letters)*[ (no. of lowercase letters)^10 ]*(no. of numbers)

= (26)*[ (26)^10 ]*10

= 3.6703445 * 10^16

no. of passwords = 3.6703445 * 10^16

time to test = (no. of possible passwords) * (time to test each)

= ( 3.6703445 * 10^16 ) * (1 nanosecond)

= 3.6703445 * 10^16 nanoseconds

c.

Passwords that start with a capital letter, have between 10 and 20 lowercase letters,and end with a number.

no. of passwords

= (no. of uppercase letters)*[ (no. of lowercase letters)^10 ]*(no. of numbers) + (no. of uppercase letters)*[ (no. of lowercase letters)^11 ]*(no. of numbers) + ...... + (no. of uppercase letters)*[ (no. of lowercase letters)^20 ]*(no. of numbers)

= (26)*[ (26)^10 ]*10 + (26)*[ (26)^11 ]*10 + (26)*[ (26)^12 ]*10 + (26)*[ (26)^13 ]*10 + (26)*[ (26)^14 ]*10 + (26)*[ (26)^15 ]*10 + (26)*[ (26)^16 ]*10 + (26)*[ (26)^17 ]*10 + (26)*[ (26)^18 ]*10 + (26)*[ (26)^19 ]*10 + (26)*[ (26)^20 ]*10

= 5.3885715 * 10^30

time to test = (no. of possible passwords) * (time to test each)5.3885715* 10^30 * 1 nanosecond

= 5.3885715 * 10^30 nanoseconds

d.

for each unit of the password there are 20,000 options

no. of possible passwords = (20,000)^4

= 1.6 * 10^17

time to test = (no. of possible passwords) * (time to test each)

= (1.6 * 10^17 ) * (1 nanosecond)

= 1.6 * 10^17 nanoseconds

(please UPVOTE)